OpenStudy (anonymous):

please i need help ASAp !!!!! ok so i have a rectangle with the base = 5/4h+10 and height= h the teacher asks for a an expression for the area Area=5/4h^2+10h now i have to do it for a triangle with Height= 2/5y+4 and base=1/2y pleaseee help its due right at 11 pm

OpenStudy (gw2011):

Area of a triangle=(1/2)(b)(h) A=(1/2)(1/2y)(2/5y+4) A=(1/4y)(2/5y+4) A=1/10y^2+y